OpenStudy (anonymous):

Can someone please help When dilute sulfuric acid is added to magnesium, 0.280 liter of hydrogen gas, at S.T.P., are produced. find the: a) mass of magnesium consumed. b) mass of magnesium sulfate produced. THANK YOU!!

OpenStudy (chmvijay):

can u write the balanced equation for the following

OpenStudy (anonymous):

Mg (s) + H2SO4 (aq) -> MgSO4 (aq) + H2 (g)

OpenStudy (anonymous):

n(H2)=n(MgSO4)=n(Mg)=0.28/22.4= m(Mg)=n(Mg)*24= m(MgSO4)=n(MgSO4)*(24+96)=

OpenStudy (anonymous):

i've done it this way too thank you :)